# Mean-field limits for Consensus-Based Optimization and Sampling

### v3

@@ A.3 Generalized Wasserstein stability estimates
  • for all $R>0$ and $p>1$, there is $C=C(p,R)$
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant C\,W_{p}(\mu,\nu),\qquad\varrho:=\frac{(s+1)p}{p-1}$ (A.13)
  • for all $R>0$ and $p>1$, there is $C=C(p,R)$
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\rvert\leqslant C\,W_{p}(\mu,\nu),\qquad\varrho:=\frac{(s+2)p}{p-1}$ (A.14)
@@ Acknowledgments
We are grateful to Yunan Yang, Arnaud Guyader and Louis-Pierre Chaintron for useful discussions, and to the Lorentz Center (Leiden, the Netherlands) for hosting the workshop “Purpose-driven particle systems” in March 2023, which enabled discussions about this work. FH is supported by start-up funds at the California Institute of Technology. FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM.

### v4

@@ A.3 Generalized Wasserstein stability estimates
  • for all $R>0$ and $p>1$, there is $C=C(p,R)$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant C\,\mathcal{W}_{p}(\mu,\nu),\qquad\varrho\coloneq\frac{(s+1)p}{p-1}$ (A.13)
  • for all $R>0$ and $p>1$, there is $C=C(p,R)$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\rvert\leqslant C\,\mathcal{W}_{p}(\mu,\nu),\qquad\varrho\coloneq\frac{(s+2)p}{p-1}$ (A.14)
@@ Acknowledgments
We are grateful to Yunan Yang, Arnaud Guyader and Louis-Pierre Chaintron for useful discussions, to Hui Huang and Hicham Kouhkouh for pointing out an issue with Assumption (H2), and to the anonymous reviewers for very useful comments and suggestions. We also thank the Lorentz Center (Leiden, the Netherlands) for hosting the workshop “Purpose-driven particle systems” in March 2023, which enabled discussions about this work. FH is supported by start-up funds at the California Institute of Technology and by NSF CAREER Award 2340762. FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM. UV is partially supported by the European Research Council (ERC) under the EU Horizon 2020 programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grants ANR-21-CE40-0006 (SINEQ) and ANR-23-CE40-0027 (IPSO).